a chemist needs 120 milliliters of a 50% solution but has only 48% and 60% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
Answer by stanbon(75887)

a chemist needs 120 milliliters of a 50% solution but has only 48% and 60% solutions available. Find how many milliliters of each that should be mixed to get the desired solution.
-----------------
Equations:
Quantity: x + y = 120 ml
Active::: 0.48x + 0.60y = 0.50*120
------------------
Multiply thru 1st by 48
Multiply thru 2nd by 100
---------------------------
48x + 48y = 48*120
48x + 60y = 50*120
=======================
12y = 2*120
y = 20 ml (amt. of 60% solution needed in the mix)
---
Since x + y = 120 , x = 100 ml (amt. of 48% solution needed in the mix)
===================
Cheers,
Stan H.
====================